# Connection between gravity and quantum mechanics: an algebraic approach to cosmological issues

### v1

@@ VIII Planck-scale physics, the hot Big Bang, and the Universe’s accelerating expansion
If we traverse backward through the ’arrow of time’, we found spacetime,which is fundamentally three-dimensional and granular; interactions within this framework may lead to photons arranging themselves into a two-dimensional condensate as a lower-energy or stable form. This implies that the 3D spacetime ”background” dictates the conditions that allow the 2D condensate to form and persist. These condensed photons having substantial Plank energy and mass densities, are contained within a distinct, non-Euclidean spacetime, characterized by a circumference-to-radius ratio of one. This unique geometry emerges from the substantial Planck gravitational potential, $\phi_{p}$, which is equivalent to $c^{2}$, thus confirming the peculiar structure of this domain. Notwithstanding its granular and highly energetic characteristics, this 3D Planck photon condensate establishes an exceptionally dense, spherical arrangement. In a broader perspective, this dense condensate functions as a singular entity with a conventional Euclidean metric, integrating the non-Euclidean characteristics of individual photons into a unified, macroscopic structure.
@@
In this first stage, we take the universe as a “granular” spacetime structure that is compact, uniform, and “granular” [68, 69, 70, 71]. Each unit is made up of Planck-scale photons that are packed together into spherical, high-energy packets. Each Planck-condensed photon in this tightly packed, granular spacetime has twelve other photons surrounding it in a very symmetric arrangement. As an example, we can think of the FCC structure of atoms/particles, which is the most dense configuration having twelve nearest neighbors in 3D.
@@
This equation serves to describe the likelihood of certain structural transformations occurring in the aftermath of the explosion, such as the production of photon pairs, chains, or clusters, which might subsequently spread and contribute to the emergent structure of spacetime and matter in the early universe. Solving this twelfth-degree equation would provide a statistical foundation for understanding the complicated dynamics driving the early microstates and their ability to create larger structures or liberate photons in the immediate aftermath of the Big Bang.
@@
    the average energy density of free Plank light is
@@
It is also possible to compute the duration of the process in two different methods. The wavelength of a Planck-free photon is $2\pi$ times longer than that of a Planck condensed photon, which is equivalent to the Planck length. Naturally, the implementation of such a wavelength necessitates a duration equivalent to the $2\pi t_{p}$. The second analogous solution arises from the epoch of the formation of Planck black holes. According to the principle of causality, it is equivalent to half the length of its geodesic line divided by the speed of light, specifically $2\pi t_{p}$. This apparent alignment of results is likely, not coincidental; rather, it reflects a profound synchronization inherent to all processes at the universe’s inception, indicative of the fundamental symmetries and coherence present at the time of the Big Bang.
The cosmological constant ($\Lambda$) is a primary parameter in explaining the universe’s accelerated expansion. Calculations indicate that $\Lambda$ could theoretically reach a maximum value around $\frac{1}{l_{p}^{2}}$, where $l_{p}$ is the Planck length. This theoretical peak, about $10^{122}$ times greater than today’s observed value, would correspond to early-universe conditions of intense energy density and rapid expansion. This highest value of $\Lambda$ matches a Hubble parameter on the order of $\frac{1}{t_{p}}$, which is the Planck time. This means that the universe is expanding as quickly as it can under the limits of quantum gravity [73, 74, 75, 76].
The previous “vacuum energy” problem needs to be fixed, which said that the cosmological constant was about $10^{120}$ times bigger than what was observed. This is what has been called the “cosmological constant problem” or “vacuum catastrophe,” which is one of the biggest differences between theory and observation in physics [77, 78, 79].
By proposing that $\Lambda$ reaches its maximum value only under early-universe conditions (near the Planck scale), this approach may nullify the enormous factor of $10^{120}$ and indicates that the observed, much smaller value of $\Lambda$ reflects a natural evolution from this maximum rather than a fine-tuning issue. This revised perspective potentially resolves the cosmological constant problem by aligning $\Lambda$ with realistic early-universe scenarios, where its value would have been vastly higher than it is now.
@@
    With a probability $\frac{1-\aleph}{2}-\aleph=\frac{1-3\aleph}{2}$ the plank blackhole decay occures:
@@
The second result indicates the complete extinction of PBHs that emitted significant light energy, thereby enabling the formation of the next generation. Prior research concerning the principles of quantum absorption indicates that, in the context of radiative transfer between black holes, the mass of the latter must be not less than twice that of the former.
However, what is the source of the final entities? To this point, our analysis has been confined to average, representative values concerning the formation of black holes. In accordance with the foundational concepts of probability theory, it is essential to examine a comparatively small yet impactful quantity of black holes capable of producing a new generation characterized by mass/energy density,
@@
IX The growth of supermassive black holes and very early clumping of the universe
@@
As established earlier, successive generations of black holes formed primarily through the radiative transfer of light energy. This process involved energy transfer from one generation, which was significantly more numerous but “lighter,” to a subsequent generation that was far less numerous but “heavier.” This transfer of energy allowed for the growth and evolution of PBHs over time, leading to the formation of SMBHs at the centers of galaxies. These SMBHs play a crucial role in shaping the structure and dynamics of galaxies. This can be articulated straightforwardly and lucidly with significant implications,
@@
Based on the fact that both the quantum information and the growing black hole’s accretion area are proportional to the quantum number, we can write a growth equation for $n_{s}(t)$ in accordance with Bohr’s correspondence principle as:
@@
It should be noted that the time-dependent mass function can also be found in dynamical spacetimes like those of the Vaidya spacetime [87, 88, 89], and aligns with models involving time-dependent mass functions for SMBHs [90].
@@
The growth of SMBHs occurs through a cyclical mechanism in which successive generations of PBHs generate significant photon emissions that facilitate the growth of subsequent black holes. Fundamental Eqs. (15) and (24) describe this cyclic mechanism, which results in significant increases in luminosity with each cycle while the SMBH’s accretion area expands at an even faster rate, enhancing its gravitational pull. The gravitational influence promotes the aggregation of PBHs, which signify the evolution of dark matter in the early universe. The duration (23) of each subsequent cycle increases significantly, thereby improving the formation efficiency of early galactic precursors and SMBHs of different sizes. As these cycles end, the universe cools, leading to a significant reduction in the growth of SMBHs and signaling a shift to a less active cosmological phase.
@@
X Baryogenesis through primordial black hole production
The most convincing evidence of our model is that PBHs are directly tied to baryogenesis. It is well-known that dark matter surrounds baryonic matter everywhere [93, 94, 95]. Moreover, we find, in general, that where the concentration of dark matter is higher on large scales, there is a corresponding increase in the concentration of ordinary matter.
@@ X Baryogenesis through primordial black hole production
Our focus lies in establishing a correlation between this imbalance and the mechanisms underlying particle pair formation by black holes, utilizing a systematic approach. This framework makes it possible to find strong links between asymmetry in the real world and mechanisms that work in theory. This helps us learn more about how particles form in the early universe and when gravity is very strong.
@@
where $E(n^{0})=0.9396GeV$ is the rest energy of the neutron, $E(c_{q})=1.25GeV$ energy of the free “charm” quark, next in this parameter after neutron. This constraint implies that PBHs predominantly generate neutrons up to a specific quantum energy level $n_{s}$. The highest limit, $E(c_{q})$, ensures that the PBHs won’t have enough energy to produce charm quarks, restricting particle production to neutrons and lighter particles. By limiting $n_{s}$ to the specified range, the model ensures that the PBH decays predominantly yield baryons, such as protons, rather than other particles that would not enhance baryon asymmetry. The concept assumes that gravitational effects or primordial black hole decay can induce CP violation, essential for generating a matter-antimatter imbalance. PBHs create conditions in which gravitational interactions may inherently result in such violations, eliminating the necessity for alternative CP-violating mechanisms. Here we need the “charm” quark rest energy in order to determine the quantum number $n_{s}(c_{q})$ at which neutron production stops. Using a simple equation
$\frac{[2n^{*}_{s}(n^{0})-2n_{s}(c_{q})]E(n^{0}){\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}]}}{E_{p}\sqrt{n_{s}^{*}(n^{0})}}=\frac{\Omega_{b}}{\Omega_{DM}}$ (78)
@@
The expression on the left side of Eq. (78) denotes the productivity of one typical black hole, expressed as the difference between the number of neutrons and the number of annihilated charm quarks it produces (square brackets) multiplied by the rest of the energy of the neutron and divided by the rest energy of typical producing black hole. We equate this to the energy density of baryon $\Omega_{b}$=0.0496 divided by the energy density of the producing black hole $\Omega_{DM}$=0.32228 (Eq. (59)). In essence, this equation as a whole expresses a delicate balance between the productive capacity of PBHs in generating baryonic matter and the observed ratio of baryonic to dark matter in the universe.
@@
According to Eq. (23) the formation of baryonic matter takes place over roughly $87$ thousand years, while the complete cycle of primordial black hole evaporation and subsequent extinction lasts approximately $190$ thousand years. During the intermediate period, from $87$ to $190$ thousand years, PBHs may continue decaying and emitting radiation, even though baryonic matter production has largely ceased. This radiation increases the universe’s overall radiation density but does not significantly affect the baryon-to-dark matter ratio, which was established during the initial baryonic production phase [102].
@@
The most important observation is that the processes involved in the generation of baryonic matter, the evaporation of black holes, and the reheating phase occurred in localized “clumps” rather than in a uniform manner. This non-uniformity played a significant role in the subsequent formation of the first galaxies, suggesting that before the phase of ordinary matter creation, the universe likely displayed inhomogeneities. The quantum perturbation [22] could also give rise to these inhomogeneities. Primordial density fluctuations can lead to the formation of gravitational wells that attract matter, thereby increasing particle interactions, promoting the formation of black holes, and resulting in the production of matter within those regions. Consequently, primordial galaxies emerged in proximity to these dense regions, where primordial black holes facilitated baryonic synthesis and localized reheating occurred, thus establishing the foundation for the large-scale structure of the Universe.
XI Formation of Modern Dark Matter and Large-Scale Cosmic Structure
@@
The hard X-ray spectrum emitted by PBHs offers valuable insights into the nature of dark matter and the mechanisms potentially driving cosmic acceleration. According to Eqs. (31), (80), (81) this emission spectrum is characterized by a peak intensity at photon energies near $E_{\gamma}=30\leavevmode\nobreak\ KeV$ and typical intensity of $E_{\gamma}=40\leavevmode\nobreak\ KeV$, with an upper cutoff observed at around $250\leavevmode\nobreak\ KeV$. This exactly matches the unresolved origin of the cosmic hard X-ray background (CXR), which was discovered before the famous cosmic microwave background (CMB). The spectral properties are important as they may clarify the mass distribution of PBHs and impose constraints on dark matter theories, considering the proposed role of PBHs as candidates for dark matter [11, 23]. Furthermore, the X-ray emissions contribute to our understanding of high-energy processes in the early universe, influencing theories on the acceleration of cosmic expansion.
@@
This distinctive hard X-ray spectrum has been observed in high-energy environments across the cosmos, including active galactic centers, galaxy clusters, and superclusters [105]. The extensive emissions indicate the potential existence of a substantial population of PBHs, supporting the theory that PBHs may represent a considerable fraction of dark matter [11, 106]. PBHs may crash into and merge in the hot, high-energy parts of SMBH coronas because of the strong gravitational fields around them. This can release a lot of X-rays. The resulting intense X-ray radiation, especially in non-jet-related phenomena, aligns with the theory that PBH activity may account for a considerable portion of the unresolved cosmic X-ray background [100, 107].
@@
Thus, we may conclude that the hard X-ray spectrum produced by PBHs offers essential information regarding the characteristics of dark matter, and the mechanisms responsible for X-ray emission in regions of high density, which also play a role in the generation of dark energy. These interactions enhance our comprehension of PBHs as candidates for dark matter and indicate their potential importance in the universe’s accelerating expansion, illustrating the relationship between dark matter and dark energy in cosmic evolution. The transition of dark matter into dark energy may possibly resolve the “core-cusp problem” [108].
@@
The JWST has revealed unexpectedly mature galaxies in the early universe, which may be explained by the rapid initial clustering of dark matter halos. This clustering generated significant gravitational potentials that drew ordinary matter into dense regions, accelerating early galaxy formation [109]. The fast growth of these dark matter halos, which create gravitational wells, would naturally speed up the formation of the universe’s structure by directing normal matter into areas with lots of it, where galaxies and stars may come together faster than was thought before. Consequently, these well-structured galaxies, identified in what was previously considered a primordial and underdeveloped epoch, exemplify the dynamic efficacy of the universe’s formative mechanisms.
@@
To the best of our knowledge we can say that black holes with masses around $10^{18}$ g (Eq.(80) have lifetimes comparable to the age of the Universe. Heavier black holes can be considered stable, while lighter ones are actively evaporating or may have already fully evaporated since their formation. Black holes with masses near $10^{18}$ g, however, could still persist today and potentially contribute to dark matter. It is clear from the above statement that the average number density of dark matter particles (or PBHs) is about $10^{41}$ times lower than that of normal matter when we look at the current mass parameter $M_{s}(\sim 10^{18}g)$ along with the known values of $\Omega_{b}$ and $\Omega_{DM}$. This is shown by the formula $\frac{\Omega_{DM}m_{p}}{\Omega_{b}M_{s}}$, where $m_{p}=1.67\times 10^{-24}g$ is the mass of a proton. This aligns perfectly with the cosmic phenomenon of the “Bullet Cluster” [110], where vast clouds of dark matter move through each other without obstruction, demonstrating their minimal interaction. In contrast, intergalactic gas clouds, which possess a notably higher density and are rich in ordinary matter, experience intense electromagnetic interactions that lead to their deceleration and dispersion when they collide. This notable disparity underscores the intriguing properties of dark matter, which interacts exclusively through gravitational forces.

### v2

@@ VIII Planck Scale to Expansion: The Quantum Foundations of Cosmology
If we traverse backward through the ’arrow of time’, we find spacetime, which is fundamentally three-dimensional and granular; interactions within this framework may lead to photons arranging themselves into a two-dimensional condensate as a lower-energy or stable form. This implies that the 3D spacetime ”background” dictates the conditions that allow the 2D condensate to form and persist. These condensed photons having substantial Planck energy and mass densities, are contained within a distinct, non-Euclidean spacetime, characterized by a circumference-to-radius ratio of one. This unique geometry emerges from the substantial Planck gravitational potential, $\phi_{p}$, which is equivalent to $c^{2}$, thus confirming the peculiar structure of this domain. Notwithstanding its granular and highly energetic characteristics, this 3D Planck photon condensate establishes an exceptionally dense, spherical arrangement. In a broader perspective, this dense condensate functions as a singular entity with a conventional Euclidean metric, integrating the non-Euclidean characteristics of individual photons into a unified, macroscopic structure.
@@
In this first stage, we take the universe as a “granular” spacetime structure that is compact, uniform, and “granular” [63, 64, 65, 66]. Each unit is made up of Planck-scale photons that are packed together into spherical, high-energy packets. Each Planck-condensed photon in this tightly packed, granular spacetime has twelve other photons surrounding it in a very symmetric arrangement. Similar to what we found in FCC structure of atoms/particles, which is the most dense configuration having twelve nearest neighbors in 3D.
@@
Where $\aleph$ is the probability of converting a condensed Planck photon to a free one and $(1-\aleph)^{12}$ is the probability for all its twelve surrounding photons to become bound, i.e. to take part at events of the type Eq. (45). This equation (42) serves to describe the likelihood of certain structural transformations occurring in the aftermath of the explosion, such as the production of photon pairs, chains, or clusters, which might subsequently spread and contribute to the emergent structure of spacetime and matter in the early universe. Solving this twelfth-degree equation would provide a statistical foundation for understanding the complicated dynamics driving the early microstates and their ability to create larger structures or liberate photons in the immediate aftermath of the Big Bang.
@@
    the average energy density of free Planck light is
@@
It is also possible to compute the duration of the process in two different methods. The wavelength of a Planck-free photon (29) is twice that of a Planck condensed photon (1), which is equivalent to the Planck length. Naturally, the implementation of such a wavelength necessitates a duration equivalent to $2\pi t_{p}$ (25). The second analogous solution arises from the epoch of the formation of Planck black holes. According to the principle of causality, it is equivalent to half the length of its geodesic line divided by the speed of light, specifically $2\pi t_{p}$ (25). This apparent alignment of results is likely, not coincidental; rather, it reflects a profound synchronization inherent to all processes at the universe’s inception, indicative of the fundamental symmetries and coherence present at the time of the Big Bang.
The cosmological constant ($\Lambda$) is a primary parameter in explaining the universe’s accelerated expansion. Calculations indicate that $\Lambda$ could theoretically reach a maximum value around $\frac{1}{l_{p}^{2}}\simeq 10^{70}cm^{-2}$, where $l_{p}$ is the Planck length. This theoretical peak, about $10^{122}$ times greater than today’s observed value, would correspond to early-universe conditions of intense energy density and rapid expansion. This highest value of $\Lambda$ matches a Hubble parameter on the order of $\frac{1}{t_{p}}$, which is the Planck time. This means that the universe is expanding as quickly as it can under the limits of quantum gravity [68, 69, 70, 71].
The so-called “vacuum energy” problem needs to be fixed, which said that the cosmological constant was about $10^{120}$ times bigger than what was observed. This is what has been called the “cosmological constant problem” or “vacuum catastrophe,” which is one of the biggest differences between theory and observation in physics [72, 73, 74].
By proposing that $\Lambda$ reaches its maximum value only under early-universe conditions (near the Planck scale), this approach may nullify the enormous factor of $10^{120}$ and indicates that the observed, much smaller value of $\Lambda$ reflects a natural evolution from its maximum rather than a fine-tuning issue. This revised perspective potentially resolves the cosmological constant problem by aligning $\Lambda$ with realistic early-universe scenarios, where its value would have been vastly higher than it is now.
@@
    With a probability $\frac{1-\aleph}{2}-\aleph=\frac{1-3\aleph}{2}$ the Planck blackhole decay occures:
@@
The second result indicates the complete extinction of PBHs that emitted significant light energy, thereby enabling the formation of the next generation. Prior research concerning the principles of quantum absorption indicates that, in the context of radiative transfer between PBH, the mass of the latter must be not less than twice that of the former.
However, what is the source of the final entities? To this point, our analysis has been confined to average, representative values concerning the formation of black holes. In accordance with the foundational concepts of probability theory, it is essential to examine a comparatively small yet impactful quantity of PBHs capable of producing a new generation characterized by mass/energy density,
@@
It should be noted that, during the Planck era, the thermodynamics of the universe was dictated by the principles of black hole thermodynamics. Contrary to the expectations of a conventional cosmological model, the universe’s primordial regions were not causally isolated; instead, it exhibited characteristics akin to black holes, that facilitate thermal equilibrium over all areas. This relationship between black hole thermodynamics and the early cosmos potentially eliminates the necessity for special explanations on how distant regions could possess identical temperatures and characteristics. Thus it may remove the so-called ’horizon problem’.
IX Formation of Supermassive Black Holes and the Universe’s Primordial Clumping
@@
As established earlier, successive generations of PBHs formed primarily through the radiative transfer of light energy. This process involved energy transfer from one generation, which was significantly more numerous but “lighter,” to a subsequent generation that was far less numerous but “heavier.” This transfer of energy allowed for the growth and evolution of PBHs over time, leading to the formation of SMBHs at the centers of galaxies. These SMBHs play a crucial role in shaping the structure and dynamics of galaxies. This can be articulated straightforwardly and lucidly with significant implications,
@@
Based on the fact that both the quantum information and the growing PBH’s accretion area are proportional to the quantum number, we can write a growth equation for $n_{s}(t)$ in accordance with Bohr’s correspondence principle as:
@@
It should be noted that the presented exponential time-dependent mass function may explain the early existence of SMBHs and the most powerful quasars detected by modern telescopes. Again, note that the time-dependent mass function can also be found in dynamical spacetimes like those of the Vaidya spacetime [81, 82, 83], and aligns with models involving time-dependent mass functions for SMBHs [84]. However, we are not discussing pure gravitational spacetime here.
@@
The growth of SMBHs occurs through a cyclical mechanism in which successive generations of PBHs generate significant photon emissions that facilitate the growth of subsequent black holes. Fundamental Eqs. (16) and (27) describe this cyclic mechanism, which results in significant increases in luminosity with each cycle while the SMBH’s accretion area expands at an even faster rate, enhancing its gravitational pull. The gravitational bonding promoted by the growth of SMBHs, acting as gravitational attractors, aggregates the clumps of PBHs into complex structures, which signify the early evolution of dark matter in the early universe. The duration (26) of each subsequent cycle increases significantly, thereby improving the formation efficiency of early galactic precursors and SMBHs of different sizes. As these cycles end, the universe cools, leading to a significant reduction in the growth of SMBHs and signaling a shift to a less active cosmological phase.
@@
X Baryogenesis driven by PBH Evolution
We intend to provide the most compelling evidence for our model, illustrating that PBHs are directly associated with baryogenesis. It is well-known that dark matter surrounds baryonic matter everywhere [87, 88, 89]. Furthermore, it turns out that regions with higher concentrations of dark matter on large scales, correspondingly exhibit increased concentrations of ordinary matter.
@@ X Baryogenesis driven by PBH Evolution
Our focus lies in establishing a correlation between this imbalance and the mechanisms underlying particle pair formation by PBHs, utilizing a systematic approach. This framework makes it possible to find strong links between asymmetry in the real world and mechanisms that work in theory. This helps us learn more about how particles form in the early universe and when gravity is very strong.
@@
where $E(n^{0})=0.9396GeV$ is the rest energy of the neutron, $E(c_{q})=1.25GeV$ energy of the free “charm” quark, next in this line after neutron. This constraint implies that PBHs predominantly generate neutrons up to a specific quantum energy level $n_{s}$. The highest limit, $E(c_{q})$, ensures that the PBHs won’t have enough energy to produce charm quarks, restricting particle production to neutrons and lighter particles. By limiting $n_{s}$ to the specified range, the model ensures that the PBH decays predominantly yield baryons, such as protons, rather than other particles that would not enhance baryon asymmetry. The concept assumes that gravitational effects or PBH decay can induce CP violation, essential for generating a matter-antimatter imbalance. PBHs create conditions in which gravitational interactions may inherently result in such violations, eliminating the necessity for alternative CP-violating mechanisms. Here we need the “charm” quark rest energy in order to determine the quantum number $n_{s}(c_{q})$ at which neutron production stops. Using a simple equation
$\frac{[2n^{*}_{s}(n^{0})-2n_{s}(c_{q})]E(n^{0})}{E_{p}\sqrt{n_{s}^{*}(n^{0})}}=\frac{\Omega_{b}}{\Omega_{DM}}$ (82)
@@
The expression on the left side of Eq. (82) denotes the productivity of one typical PBH, expressed as the difference between the number of neutrons and the number of annihilated charm quarks it produces (square brackets) multiplied by the rest of the energy of the neutron and divided by the rest energy of typical producing black hole. We equate this to the energy density of baryon $\Omega_{b}$=0.0496 divided by the energy density of the producing dark matter $\Omega_{DM}$=0.32228 (Eq. (63)). In essence, this equation as a whole expresses a delicate balance between the productive capacity of PBHs in generating baryonic matter and the observed ratio of baryonic to dark matter in the universe.
@@
According to Eq. (26) the formation of baryonic matter takes place over roughly $87$ thousand years, while the complete cycle of PBH evaporation and subsequent extinction lasts approximately $190$ thousand years. During the intermediate period, from $87$ to $190$ thousand years, PBHs may continue decaying and emitting radiation, even though baryonic matter production has largely ceased. This radiation increases the universe’s overall radiation density but does not significantly affect the baryon-to-dark matter ratio, which was established during the initial baryonic production phase [95].
@@
The most important observation is that the processes involved in the generation of baryonic matter, the evaporation of PBHs, and the reheating phase occurred in localized “clumps” rather than in a uniform manner. This non-uniformity played a significant role in the subsequent formation of the first galaxies, suggesting that before the phase of ordinary matter creation, the universe likely displayed inhomogeneities. The quantum perturbation [93] could also give rise to these inhomogeneities. Primordial density fluctuations can lead to the formation of gravitational wells that attract matter, thereby increasing particle interactions, promoting the formation of PBHs, and resulting in the production of matter within those regions. Consequently, primordial galaxies emerged in proximity to these dense regions, where PBHs facilitated baryonic synthesis and localized reheating occurred, thus establishing the foundation for the large-scale structure of the Universe.
XI The Evolution of Dark Matter and Its Impact on Cosmic Structure
@@
The hard X-ray spectrum emitted by PBHs offers valuable insights into the nature of dark matter and the mechanisms potentially driving cosmic acceleration. According to Eqs. (35), (84), (85) this emission spectrum is characterized by a peak intensity at photon energies near $E_{\gamma}=30\leavevmode\nobreak\ KeV$ and typical intensity of $E_{\gamma}=40\leavevmode\nobreak\ KeV$, with an upper cutoff observed at around $250\leavevmode\nobreak\ KeV$. This exactly matches the unresolved origin of the cosmic hard X-ray background (CXRB) [98, 99], which was discovered before the famous cosmic microwave background (CMB). The spectral properties are important as they may clarify the mass distribution of PBHs and impose constraints on dark matter theories, considering the proposed role of PBHs as candidates for dark matter [5, 1]. Furthermore, the X-ray emissions contribute to our understanding of high-energy processes in the early universe, influencing theories on the acceleration of cosmic expansion.
@@
This distinctive hard X-ray spectrum has been observed in high-energy environments across the cosmos, including active galactic centers, galaxy clusters, and superclusters [98, 99, 100]. The extensive emissions indicate the potential existence of a substantial population of PBHs, supporting the theory that PBHs may represent a considerable fraction of dark matter [5, 101]. PBHs may crash into and merge in the hot, high-energy parts of SMBH coronas because of the strong gravitational fields around them. This can release a lot of X-rays. The resulting intense X-ray radiation, especially in non-jet-related phenomena, aligns with the theory that PBH activity may account for a considerable portion of the unresolved cosmic X-ray background [102, 103, 104].
@@
Thus, we may conclude that the hard X-ray spectrum produced by PBHs offers essential information regarding the characteristics of dark matter, and the mechanisms responsible for X-ray emission in regions of high density, which also play a role in the generation of dark energy. These interactions enhance our comprehension of PBHs as candidates for dark matter and indicate their potential importance in the universe’s accelerating expansion, illustrating the relationship between dark matter and dark energy in cosmic evolution. This transition of dark matter into dark energy may possibly resolve the “core-cusp problem” [105] which refers to a discrepancy between the predicted and observed density profiles of dark matter in the centers of galaxies, particularly in galaxy clusters and dwarf galaxies.
@@
The JWST has revealed unexpectedly mature galaxies in the early universe, that strongly support the above-presented results, and, which may not be explained by the arguments, presented in [106]. The fast growth of dark matter halos, which create gravitational wells, would naturally speed up the formation of the universe’s structure by directing normal matter into areas with lots of it, where galaxies and stars may come together faster than what was thought before. Consequently, these well-structured galaxies, identified in what was previously considered a primordial and underdeveloped epoch, exemplify the dynamic efficacy of the universe’s formative mechanisms.
To the best of our knowledge, we can say that black holes with masses around $10^{18}$ g (Eq.(81)) have lifetimes much more than the age of the Universe (at about 50 million times more!).
@@
It is clear from the above statement that the average number density of dark matter particles (or PBHs) is about $10^{41}$ times lower than that of normal matter when we look at the current mass parameter $M_{s}(\sim 10^{18}g)$ along with the known values of $\Omega_{b}$ and $\Omega_{DM}$. This can be calculated by the formula $\frac{\Omega_{DM}m_{p}}{\Omega_{b}M_{s}}$, where $m_{p^{+}}=1.67\times 10^{-24}g$ is the mass of a proton. This aligns perfectly with the cosmic phenomenon of the “Bullet Cluster” [107], where vast clouds of dark matter move through each other without obstruction, demonstrating their minimal interaction. In contrast, intergalactic gas clouds, which possess a notably higher density and are rich in ordinary matter, experience intense electromagnetic interactions that lead to their deceleration and dispersion when they collide. This notable disparity underscores the intriguing properties of dark matter, which interacts exclusively through gravitational forces.